x^3- 3x^2-4x+12=0
i have to solve for x

Answer by checkley75(3666)

x^3-3x^2-4x+12=0
x^2(x-3)-4(x-3)=0
(x^2-4)(x-3)=0
(x+2)(x-2)(x-3)=0
x+2=0
x=-2 answer.
x-2=0
x=2 answer.
x-3=0
x=3 answer.